### v1

# Stability of dark solitons in a bubble Bose-Einstein condensate

Abstract
The dynamic stability of dark solitons trapped on the surface of a two-dimensional spherical bubble is investigated. In this spherical geometry of the Bose-Einstein condensate, dark solitons are found to be unstable for the interaction parameter $\varepsilon\gtrsim 8.37$, since discrete angular modes drive snake instabilities, with the generation of vortex dipoles. We show analytically and numerically that, for each angular mode $m\geq 2$, there exists exactly one unstable mode whose dominance determines the number $m$ of vortex dipoles. Time-dependent simulations confirm the formation of vortex dipoles.
Bose-Einstein condensed gases in spherical geometries have recently attracted attention due to experiments performed aboard the International Space Station in a microgravity environment with ultra-cold gases confined in spherical and/or ellipsoidal surfaces [1, 2]. These experimental investigations were designed based on previous theoretical work on shell-like potentials [3, 4, 5]. Ground-based experiments also revealed shell bubbles, by exploring two species [6], or intending to observe alternative two-dimensional (2D) closed geometries [7]. Such investigations have shed new light on the physics of low-dimensional quantum gases, especially in closed 2D shells. See Ref. [8] for a recent overview of the present status and perspectives on quantum gases in bubble traps. Several interesting problems have been studied in this context, considering fundamental physics properties in shell-like structures [9, 10, 11, 12, 13, 14], vortex dynamics and stability [16, 15, 17], dipole interaction effects [18, 19], Berezinskii-Kosterlitz-Thouless (BKT) transition [20], thermodynamic properties of the gas adiabatic expansion from filled sphere to hollow one [21], and to our main interest, some attention has been paid to the properties of condensate mixtures trapped on a bubble [23, 22, 24]. Furthermore, the study of dimers on a spherical surface shows that the dimers are squeezed in the direction orthogonal to the center of mass motion, qualitatively changing its geometry, from 2D to one-dimensional (1D), leading to two-soliton motion on a bubble surface [25].
In the context of the properties of Bose-Einstein condensates (BECs) on spherical surfaces, a topic of particular interest concerns the existence, stability, and propagation of dark solitons [26]. Bubble BECs are realized when atoms are confined to a thin spherical shell, creating unique curvature effects that influence the dynamics of solitons. Dark solitons are shape-localized propagating waves manifested by depressions in the condensate density. On a bubble surface, geometric effects can alter the instability dynamics of dark solitons, as compared to their propagation on flat surfaces. The curved surface is expected to affect their dynamics, propagation speed, and stability. Unlike in flat quasi-1D or quasi-2D BECs, where dark solitons are susceptible to snake instability [27], decaying into vortices, the closed topology of a bubble is expected to alter the decay pathways, with the curvature introducing geometric constraints affecting the stability dynamics. Therefore, the soliton can experience strain due to the curvature, leading to contraction or expansion depending on the interactions.
The instability of a dark soliton stripe on a sphere follows a procedure similar to that in flat space. However, a sphere has no edges, such that the curvature influences perturbations that would be symmetric in a flat system. The most critical difference in a bubble refers to the fragmentation into vortices, as the resulting vortices are not free to move arbitrarily. The snake instabilities on a flat plane produce vortex-antivortex pairs that can move apart from each other and eventually are ejected from the high-density region. On a sphere, there is no edge to eject vortices as in flat space, with vortex-antivortex pairs remaining trapped on the surface, leading to complex dynamics like vortex-antivortex annihilation or formation of stable stationary patterns. The profound topological constraint implies that a single vortex is not allowed on a sphere, as vortices must exist in pairs of opposite circulation whose charges sum to zero.
Dark solitons in a flat BEC are known to oscillate and decay due to quantum fluctuations or dissipation. In quasi-1D traps, dark solitons can be metastable and decay due to transverse instabilities [28, 29]. Using harmonic traps, they were studied in [30, 31], both analytically and computationally, where it was found that the large-amplitude field modulations at a frequency resonant with the energy of a dark soliton give rise to a state with multiple vortices. The stability spectrum of the dark soliton contains complex frequencies, which disappear for sufficiently small numbers of atoms or for a large transverse confinement. The relationship between these complex modes and the snake instability was investigated numerically by real-time propagation, see also  [32]. In ring geometries, dark solitons in BEC systems, introduced in [33], were further investigated in [34].
Two-component BECs were considered in [35], where the dark soliton exists in one of the condensate components and the soliton nodal plane is filled with the second component. The filled solitons are stable for hundreds of milliseconds. By selectively removing the filling, one can make the soliton more susceptible to dynamical instabilities. For a condensate in a spherically symmetric potential, these instabilities cause the dark soliton to decay into stable vortex rings. By studying the oscillations and interactions of dark and dark-bright solitons, it was shown in  [36] that the stability of solitons can be controlled in BECs, via confinement.
The aim of the research we present concerns a study on the spectral stability of dark solitons under discrete $m$-angular modes within a BEC confined on the surface of a bubble in the approximation of a 2D spherical hollow shell [15]. We show the occurrence of exactly one unstable mode for each $m\geq 2$. These unstable modes induce snake-like excitations, which break the condensate into pieces, as observed in [15]. As verified numerically, an $m$-dominated unstable mode is followed by the creation of $m$ vortex-antivortex pairs, in agreement with our analytical predictions.
This letter is organized as follows. We first present the theoretical model for an atomic BEC confined to the surface of a perfect sphere, named a bubble. Dark solitons, together with the asymptotic approximations in the limit of small and large chemical potentials, are obtained analytically and numerically, supported by Supplemental Material (SM) [37]. We then report the analytical results of the corresponding stability analysis, demonstrating they are in excellent agreement with numerical observations. Finally, we report outcomes of the nonlinear dynamics of the system, from which it follows that the instabilities result in the formation of vortex dipoles on the dark soliton condensate.
Mathematical model — The present study is performed by assuming the condensate is trapped on the surface of a rigid spherical shell, aiming to mimic the cold-atom bubble experiments that are currently being performed in microgravity environments. In this approach, the system can be studied by considering a reduction of the 3D Gross-Pitaevskii equation (GPE) to a corresponding 2D system described by spherical coordinates $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$, with fixed radius $R>0$, which is also assumed to be the space unit. In the dimensional reduction, $\delta R$ is also taken as being the radial thickness of the spherical bubble. The two-dimensional approximation is reasonable since excitations in the radial direction are inaccessible due to the large amount of energy required for them. This is true when the thickness $\delta R$ of the 3D spherical shell is very small compared to its radius $R$, that is, $\delta R\ll R$, as discussed in Refs.[15, 22]. We also stress that our main concern is the dynamic stability of the dark soliton. We do not take into account its energetic instability, which could be triggered by a thermal cloud that is neglected here. The condensate can be described in the mean-field approach as a system of two coupled GPEs [38, 39] with the nonlinear two-body parameter $\overline{g}\equiv{4\pi\hbar^{2}{a}N}/{M}$, where $N$ is the number of atoms with mass $M$, and $a$ is the atom-atom $s$-wave scattering length. With this definition for the two-body contact interaction, the corresponding wave function is assumed normalized to one in the $L^{2}$-norm. By giving $R$ as the length unit, with the time in units of $MR^{2}/(2\hbar)$, the dimensionless 2D GPE for the wave function $\psi\equiv\psi(\theta,\phi,t)$ is written as
${\rm i}{\partial_{t}\psi}$ $=$ $-\Delta_{\rm 2D}\psi+g{\left|\psi\right|}^{2}\psi$
$\equiv$ $-\left[\frac{1}{\sin\theta}{\partial_{\theta}}\left(\sin\theta{\partial_{\theta}}\right)+\frac{1}{\sin^{2}\theta}{\partial^{2}_{\phi}}\right]\psi+g{\left|\psi\right|}^{2}\psi,$
where $g\equiv{4\sqrt{2\pi}aN}/{\delta R}$. With $\ell$ and $m$ being the quantum numbers related to the angular variables $\phi$ and $\theta$, a given state $\psi$ is given by $\psi_{\ell m}$. By separating the $m$-angular mode, with $\psi_{\ell m}\equiv e^{{\rm i}m\phi}\Theta_{\ell m}(\theta)$, the above defined $\Delta_{\rm 2D}$ is reduced to the $1D$ operators $\Delta_{m}$, leading to the Laplace equation
$\Delta_{m}\Theta_{\ell m}=-\left[\frac{d^{2}}{d\theta^{2}}+\cot\theta\;\frac{d}{d\theta}-\frac{m^{2}}{\sin^{2}\theta}\right]\Theta_{\ell m}=\lambda_{\ell}\Theta_{\ell m}.$ (2)
Recall that bounded solutions of this equation exist if and only if $\lambda_{\ell}=\ell(\ell+1)$, with $\ell\in\mathbb{N}_{0}=\{0,1,2,\dots\}$, and $\ell\leq m\leq\ell$. They can be expressed by the associate Legendre polynomials, with $\Theta_{\ell m}(\theta)=P_{\ell}^{m}(\theta)$.
Existence of dark solitons — The existence of dark solitons on a bubble (dark ring solitons), previously investigated in [33, 34], can be associated with the stationary solution of the GPE,
$-\left(\frac{d^{2}}{d\theta^{2}}+\cot\theta\;\frac{d}{d\theta}\right)f(\theta)$ $+$ $\varepsilon|f(\theta)|^{2}f(\theta)=\mu f(\theta),$ (3)
where $\varepsilon=g/(2\pi)$ is the magnitude of the defocusing nonlinearity, and $\mu$ is the chemical potential. To obtain (3) from (Stability of dark solitons in a bubble Bose-Einstein condensate), we assume
$\psi_{s}(\theta,\phi,t)=\frac{f(\theta)}{\sqrt{2\pi}}e^{-i\mu t},\;\;\int_{0}^{\pi}\sin{\theta}|f(\theta)|^{2}\text{d}\theta=1.$ (4)
Eq. (3) admits the conservation of flux,
$J=\sin\theta\left[\bar{f}(\theta)\frac{df(\theta)}{d\theta}-\frac{d\bar{f}(\theta)}{d\theta}f(\theta)\right],\quad 0<\theta<\pi,$ (5)
where $f(\theta)$, $df(\theta)/d\theta$ are bounded at zeros of $\sin\theta$. Hence, we have $J=0$ and, without loss of generality, we can consider a real-valued function $f(\theta):[0,\pi]\to\mathbb{R}$. The dark soliton profile is defined by a density that decreases monotonically, $|f(\theta)|^{2}$ on $0,\pi]$, with zero at $\theta=\frac{\pi}{2}$. Given the normalization (4), $\mu$ is defined by $\varepsilon$, as
$\mu=\int_{0}^{\pi}\sin{\theta}\left(\left|\frac{df(\theta)}{d\theta}\right|^{2}+\varepsilon|f(\theta)|^{4}\right)\text{d}\theta.$ (6)
A constant solution of (3) exists for every $\varepsilon>0$ in the form: $f(\theta)={1}/{\sqrt{2}}$, with $\mu=\varepsilon/{2}.$ A dark soliton is obtained by the small $\varepsilon$ limiting solution of (3):
$f(\theta)=\sqrt{\frac{3}{2}}\cos\theta+\varepsilon f_{1}(\theta)+\mathcal{O}(\varepsilon^{2}),\;\;\mu=2+\frac{9}{10}\varepsilon+\mathcal{O}(\varepsilon^{2}),$ (7)
where $f_{1}(\theta)$ is uniquely defined and proportional to $P_{3}(\cos\theta)$ [see (IS9) in the SM [37]]. As the main term of (7) is the normalized $P_{1}(\cos\theta)$, the orthogonality of the Legendre polynomials imply that the normalization (4) for small $\varepsilon$ is restricted to $\mathcal{O}(\varepsilon^{2})$.
In the other extreme, with $\varepsilon\to\infty$, the dark soliton profile becomes concentrated near the equator ($\theta=\pi/2$), according to the exact solution in the flat space:
$-\frac{d^{2}}{d\theta^{2}}f_{\infty}(\theta)$ $+$ $\varepsilon f_{\infty}^{3}(\theta)=\mu_{\infty}(\varepsilon)f_{\infty}(\theta),$ (8)
$f_{\infty}(\theta)$ $=$ $\frac{1}{\sqrt{2}}\tanh\left[\frac{\sqrt{\varepsilon}}{2}\left(\frac{\pi}{2}-\theta\right)\right],\;\mu_{\infty}(\varepsilon)=\frac{\varepsilon}{2},$
With the convenient variable change $z\equiv\frac{\sqrt{\varepsilon}}{2}\left(\frac{\pi}{2}-\theta\right)$, notice that $f_{\infty}(\theta)$ is normalized to one only if $\sqrt{\varepsilon}\to\infty$. It can be incorporated into the asymptotic expansion, as shown in [37],
$f(\theta)=f_{\infty}(\theta)+\frac{f_{\infty}^{(1)}(\theta)}{\sqrt{\varepsilon}}+\mathcal{O}\left(\frac{1}{\varepsilon}\right),\mu=\frac{\varepsilon}{2}+\sqrt{\varepsilon}+\mathcal{O}\left(1\right),$ (9)
with $f_{\infty}^{(1)}(\theta)$ being uniquely defined, given by (IS17).
In Fig. 1(a) we present profiles of the dark soliton obtained numerically, by using the shooting method described in [37]. The profile $f(\theta)$ is monotonically decreasing with the only zero at $\theta=\pi/2$ (the equator). It is close to $\sim\cos\theta$ as $\varepsilon\to 0$ according to (7) and close to $f_{\infty}(\theta)$ as $\varepsilon\to\infty$ according to (Stability of dark solitons in a bubble Bose-Einstein condensate)–(9). We confirm the asymptotic predictions by plotting $\mu$ versus $\varepsilon$ in Fig. 1(b), which shows a good agreement between the theory and numerical results.
Figure 1: (Color on-line) In (a), we show profiles of dark solitons, $f(\theta)$ (upper panel), with corresponding densities $|f(\theta)|^{2}$ (lower panel), as function of $\theta$, for three different values of $\mu$, with $\varepsilon$ obtained numerically by using the shooting method. In (b), the chemical potential $\mu$ is presented as a function of $\varepsilon$ for dark solitons. Solid lines correspond to the numerical data, with dashed and circles referring to $\mu=2+\frac{9}{10}\varepsilon$ and $\mu=\frac{\varepsilon}{2}+\sqrt{\varepsilon}+1$, respectively, in agreement with (7) and (9).
Stability of dark solitons — Linear stability analysis [40] can be performed to verify the stability of dark solitons. In this context, the well-known Bogoliubov-de Gennes (BdG) method was considered in [15]. Within this approach, small amplitude oscillations are assumed around the stationary solution (4), which is redefined as
$\psi(\theta,\phi,t)=\psi_{s}(\theta,\phi,t)+\frac{e^{-{\rm i}\mu t}}{\sqrt{2\pi}}\left[u(\theta,\phi,t)+{\rm i}v(\theta,\phi,t)\right],$ (10)
with $u$ and $v$ being real-valued. From Eqs. (Stability of dark solitons in a bubble Bose-Einstein condensate), (4), and (10), we obtain the linearized equation of motion, which leads to the following coupled equations for $u$ and $v$:
$\left\{\begin{array}[]{c}\partial_{t}u=-\Delta_{\rm 2D}v+\varepsilon f(\theta)^{2}v-\mu v,\\ -\partial_{t}v=-\Delta_{\rm 2D}u+3\varepsilon f(\theta)^{2}u-\mu u,\end{array}\right$. (13)
where the chemical potential $\mu$ is a function of $\varepsilon=g/(2\pi)$, given by (6). Performing the separation of variables for $m$-angular modes in $\phi$, with $(u,v)$ given by the eigenvectors $(\hat{u}_{m},\hat{v}_{m})$ and eigenvalue $\omega$,
$u=\hat{u}_{m}(\theta)e^{{\rm i}(m\phi+\omega t)},\quad v=\hat{v}_{m}(\theta)e^{{\rm i}(m\phi+\omega t)},$ (14)
we define the spectral stability problem. For each $m$-mode separately, we have the coupled system
$\left\{\begin{array}[]{l}\omega\hat{u}_{m}=L^{-}_{m}\hat{v}_{m},\quad L^{-}_{m}=-\Delta_{m}+\varepsilon f(\theta)^{2}-\mu,\\ \omega\hat{v}_{m}=L^{+}_{m}\hat{u}_{m},\quad L^{+}_{m}=-\Delta_{m}+3\varepsilon f(\theta)^{2}-\mu,\end{array}\right$. (15)
with $\Delta_{m}$ given in (2).
The dark soliton is spectrally stable with respect to the $m$-angular mode if the corresponding imaginary part of $\omega$ is zero, for all eigenvalues of (15). Let us consider, for each mode $m$, the corresponding $\omega_{m}$. As we know from the general theory [40], all eigenvalues $\omega_{m}$ of the spectral stability problem (15) are real if the linear operators $L^{-}_{m}$ and $L^{+}_{m}$ are positive. These linear operators enjoy two relations for applications of the comparison principle:
$L_{m}^{+}-L_{m}^{-}=2\varepsilon f^{2}(\theta)\geq 0$ (16)
and
$L_{m+1}^{\pm}-L_{m}^{\pm}=\frac{2m+1}{\sin^{2}\theta}\geq 0.$ (17)
Based on (16), in [37] we demonstrate that all eigenvalues of $L_{m}^{\pm}$ are strictly positive for $m\geq 2$ and small $\varepsilon>0$, such that all eigenvalues of $L_{m}^{+}$ remain strictly positive for $m\geq 2$ and all $\varepsilon>0$. Hence, the dark soliton is stable with respect to the $m$-angular mode for small $\varepsilon$. However, the lowest eigenvalue of $L_{m}^{-}$, defined as $\omega^{-}_{m}$, can cross zero and become negative for sufficiently large $\varepsilon$, triggering the snaking instability of the dark solitons.
Next, by considering the comparative relation (17), we demonstrate in [37] that the smallest eigenvalue of $L_{m+1}^{-}$ crosses zero for values of $\varepsilon$ larger than the smallest eigenvalue of $L_{m}^{-}$ for any given $m\geq 2$, which always occurs for large values of $m$. This implies that, for every $m\geq 2$, there exists a $\varepsilon_{m}>0$ satisfying $\varepsilon_{m}<\varepsilon_{m+1}$, such that $L_{m}^{-}$ has a simple negative eigenvalue for $\varepsilon>\varepsilon_{m}$. This further clarifies that the spectral stability problem (15) provides a real unstable eigenvalue m for $\varepsilon>\varepsilon_{m}$. As shown in Table 1, the asymptotic expression, derived in [37] for $m\gg 2$,
$\varepsilon_{m}^{th}=4m(m-1),$ (18)
provides a good approximation to predict the threshold onset of instability, considering all values of $m\geq 2$. The $\mathcal{O}(1)$ discrepancy between (18) and the numerical data represents less than 5% even for $m=2$.
Table 1: For each angular mode $m$, the threshold values [where ${\rm Im}(\omega_{m})=0$] of $\varepsilon$ are given, with $\varepsilon_{m}$ being the exact numerical results and $\varepsilon_{m}^{th}$ given by the analytical approximation (18). The respective values of $\mu$ are also presented.
$m$ 2 3 4 5 6 7
$\varepsilon_{m}$ 8.367 24.402 48.416 80.420 120.420 168.420
$\varepsilon_{m}^{th}$ 8 24 48 80 120 168
$\mu$ 8.182 18.202 32.208 50.210 72.210 98.210
The analytical predictions are confirmed numerically, as shown in Fig. 2 for the angular modes with $m=$ 2, 3, 4, and 5. The onset of instability in the spectral problem (15) corresponds to the zero eigenvalues of the linear operator $L_{m}^{-}$, with the imaginary part of the frequencies dynamically characterizing the instability. The dark soliton becomes unstable when $\varepsilon\gtrsim 8.37$, with $m=2$ being the initially dominant mode. The $m=2$ mode remains dominant for larger $\varepsilon$, until $\varepsilon\approx 35$ [when $\rm{Im}(\omega_{2})=\rm{Im}(\omega_{3})$], where the mode $m=3$ starts dominating the instability. As represented in Fig. 2, for a given $\varepsilon$, the largest imaginary eigenvalue characterizes the dominant unstable mode. If a given $\varepsilon$ corresponds to a dominant mode $m\geq 2$, then the dark soliton is affected by snake instability and tends to break into $m$ vortex-antivortex pairs.
Figure 2: (Color on-line) By varying $\varepsilon$, we show for different $m$-angular modes (indicated by the symbols and lines) that the imaginary part of the eigenvalues $\omega$ (positive defined), obtained from (15) (solid lines), start increasing from zero when the corresponding lowest eigenvalue $\omega^{-}_{m}$ of $L_{m}^{-}$ (dashed lines) becomes negative. For a given $\varepsilon$, the dominant unstable mode $m$ is provided by the largest ${\rm Im}(\omega_{m})$.
Figure 3: Dark-soliton dynamics, for given time instants $t$, with dominant instability modes $m=2$ (left panels, $\varepsilon=20$), $m=3$ (center panels, $\varepsilon=50$), and $m=4$ (right panels, $\varepsilon=100$), shown by the respective densities $|\psi|^{2}$. In 3D graphics, the darker the region, the less dense it is. Panels (d), (h), and (l) show the corresponding $|\psi(\frac{\pi}{2},\phi)|^{2}$ for the same instants $t$ represented in the upper 3D plots. The dashed lines are for $t=0$; the red dotted-dashed, when snake instabilities occur; and the solid-blue lines, after the breakup in vortex pairs.
Regarding the spectral stability of dark solitons for $m=1$, we demonstrate in [37] that the Eq. (15) has only real eigenvalues $\omega$, with no instability bifurcations. Finally, for $m=0$, it is also shown in [37] that there exists a single pair of eigenvalues $\omega$ of negative energy, which are smaller than all other pairs of eigenvalues $\omega$ of positive energy for small values of $\varepsilon$. These pairs can coalesce hypothetically for large values of $\varepsilon$, triggering another instability bifurcation. However, we have no supporting numerical evidence for the occurrence of such instability.
In Fig. 3 we illustrate the dynamical instability of the dark soliton in a sphere for three different dominant modes: $m=2$, $m=3$, and $m=4$. The three choices are taken respectively at $\varepsilon=$ 20, 50, and 100 [See Fig. 2 for the given data]. At $t=0$, we observe the dark-soliton positions at the equator, with their respective widths being smaller for larger values of $\varepsilon$. Next, in the time evolution, we can verify the onset of snake instabilities provoking the breakup of such dark solitons in vortex-antivortex pairs. As seen in panels (c) and (d), for $m=2$, two pairs of vortices are formed; in (g) and (h), for $m=3$, we observe the formation of three pairs of vortices; and, in panels (k) and (l), for $m=4$, we have four pairs of vortices. The widths of the dark solitons, as well as the healing lengths of the generated vortices (after the breakup of the snake instabilities), correspond to $\sim{1}/{\sqrt{\varepsilon}}\approx 0.22$ ($m=2$), 0.14 ($m=3$), and 0.10 ($m=4$).
The topology of the sphere implies that, on the onset of instability, the vortex number change must be +2+2, due to Poincaré-Hopf theorem, which states that any continuous tangent vector field on a sphere must have at least one point where it is zero. For a superfluid, this implies no single vortex on a sphere. They must exist in pairs of opposite circulation (vortex-antivortex pairs) whose charges sum to zero.
The numerical approach considered in the dark-soliton dynamics is carried out first by obtaining the stationary solutions using the shooting method [49, 48]. The time evolution employs the combined spectral method together with the finite difference method introduced in [15]. For the full numerical calculations of the GPE (Stability of dark solitons in a bubble Bose-Einstein condensate), we have assumed a time-step $\delta t=10^{-5}$, with spatial grids in $\theta$ and $\phi$ directions of sizes $256\times 256$. The respective step sizes were $\delta\theta=\pi/256\approx 0.0123$ and $\delta\phi=2\pi/256\approx 0.0245$. Details of the numerical methods are given in [37].
Conclusions — In the present letter, we have reported results obtained numerically and analytically for the dynamical stability of dark solitons in a single atomic BEC trapped on the surface of a rigid spherical bubble with radius $R$ and thickness $\delta R$. All the numerical and analytical results rely on a single parameter $\varepsilon$, which encompasses all information on the repulsive two-body interaction, the number of atoms, and the thickness of the bubble surface. The results of our analysis, in terms of the azimuthal excitation angular momentum modes $m$, should be useful when considering possible experimental realizations. We expect these results to impact investigations of BEC systems in spherical geometries, which are being performed aboard the international space station [2], as well as in shell condensates made with atomic mixtures [6]. Based on the spectral stability analysis, our study in the sphere shows that for the interaction parameter $\varepsilon\lesssim 8.37$, dark solitons are stable for all $m$-angular mode excitations. For $\epsilon\gtrsim 8.37$, modes with $m\geq 2$ are excited, where the dominant $m$-angular mode induces snake-like instabilities that cause the soliton to break up with the production of $m$- vortex pairs. In summary, we are highlighting the occurrence of a phenomenon that has not been previously reported when considering predictions from purely analytical studies compared with exact computational results. Our predictions for a single BEC in a sphere should provide valuable information for ongoing experimental investigations.
Acknowledgements — We thank partial support from Fundação de Amparo à Pesquisa do Estado de São Paulo [Procs. 2024/01533-7], Conselho Nacional de Desenvolvimento Científico e Tecnológico [Procs. 303263/2025-3 (LT) and 306219/2022-0 (AG)] and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (RWS).